M 378K 57470 Spring ’06 T Th 11:00-12:30 RLM 7.120

Instructor: C. Friedman RLM 11.140 471-5161 office hrs: T Th 2:30-3:30 or by appt. Web Page (homework assignments are listed): www.ma.utexas.edu/˜ friedman

Text and Syllabus: Mathematical Statistics with Applications, 6th ed., Wackerly, Mendenhall, Scheaffer, Duxbury. Syllabus: Selected sections of Chapters I-XI (possibly more; much of the early chapters is review from M362K.)

Prerequisites: M362K

Exams: We will have exams on 2/16 (class 10; Thurs), 3/30 (class 20; Thurs), 5/2 (Class 29; Tues). (No “Final Exam” planned [subject to Department approval])

* Exam grades will be scaled (if the unscaled average is ≤ 75) to some value in the 75 to 80 range. ** Makeups will be given very reluctantly.

Homework: Homework problems will be assigned each Tuesday in class and will be due the next Tuesday. A serious effort should be made on these exercises. It will probably be difficult to completely learn the course material without working problems; also exam questions may bear some resemblance to homework exercises. Additionally, all homework together will be counted as a grade equivalent to one exam grade. (Homework assignments will be listed on my web page - see above.)

Final Grades: In computing a final grade, I will count the three tests and one total homework grade equally. (So there will be four grades to average.)